OpenStudy (anonymous):

Marketing estimates that at a price of p dollars, it can sell q = 300,000 − 15,000p copies, but each copy costs $4 to make. What price will give the greatest profit?

OpenStudy (anonymous):

plz someone give mean answer

jimthompson5910 (jim_thompson5910):

What is the revenue function?

OpenStudy (anonymous):

300000 minus 15000 p

jimthompson5910 (jim_thompson5910):

that's q

jimthompson5910 (jim_thompson5910):

I want to know R

jimthompson5910 (jim_thompson5910):

R = q*p

OpenStudy (anonymous):

thats all it gave me

jimthompson5910 (jim_thompson5910):

you plug in what they give you for q

OpenStudy (anonymous):

so 4?

jimthompson5910 (jim_thompson5910):

no where it says q =

OpenStudy (anonymous):

Hercules Films is deciding on the price of the video release of its film Bride of the Son of Frankenstein. Marketing estimates that at a price of p dollars, it can sell q = 300,000 − 15,000p copies, but each copy costs $4 to make. What price will give the greatest profit? this is all it gave me

OpenStudy (anonymous):

???? is a minus

jimthompson5910 (jim_thompson5910):

yeah you already posted that

jimthompson5910 (jim_thompson5910):

you need to find R given q and p

jimthompson5910 (jim_thompson5910):

R = q*p R = ???

OpenStudy (anonymous):

240000?

jimthompson5910 (jim_thompson5910):

q = 300,000 - 15,000p

jimthompson5910 (jim_thompson5910):

R = q*p R = (300,000 - 15,000p)*p expand it out and tell me what you get

OpenStudy (anonymous):

60000

jimthompson5910 (jim_thompson5910):

no

jimthompson5910 (jim_thompson5910):

you should get a quadratic

OpenStudy (anonymous):

how do i need to find a derivative or something?

OpenStudy (anonymous):

can you give me A STEP BY STEP WAY OF ANSWERING THIS PROBLEM IM SO CONFUSED

jimthompson5910 (jim_thompson5910):

is this a calc class?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

what do you get when you distribute (x+y)z

OpenStudy (anonymous):

what does that mean?

jimthompson5910 (jim_thompson5910):

it's connected with R = (300,000 - 15,000p)*p

jimthompson5910 (jim_thompson5910):

I'm basically asking you to distribute the p through in R = (300,000 - 15,000p)*p